OpenStudy (anonymous):

factor if possible: y^6-216

OpenStudy (anonymous):

y^6-216 =[ (y^2)^3 - 6^3)] = (y^2 - 6)((y^2)^2 - 6y^2 + 6^2) = (y^2 - 6)(y^4 - 6y^2 + 36)

OpenStudy (anonymous):

i used the identity a^3 - b^3 = (a-b)(a^2 - ab + b^2)